# Color code decoder with improved scaling for correcting circuit-level noise

### v1

@@ B.2 Small-weight errors uncorrectable via the concatenated MWPM decoder
Then the subsequent MWPM on the red-only lattice causes a logical error. Similarly, the green sub-decoding fails if and only if
@@
Appendix C Additional numerical analyses
@@
Figure 14 presents separate analyses for the $\overline{Z}$- and $\overline{X}$-failure of the decoder under near-threshold circuit-level noise, which are analogous to Figs. 7(b) and (c). The long-term cross thresholds $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$ are estimated as 0.460% for the $\overline{Z}$-failure and 0.452% for the $\overline{X}$-failure, which differ by about 1.8%. Note that $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$ for the overall failure rate is about 0.456%, which is the average of these two.
@@
$p^{\star}_{\mathrm{circuit}}\approx 3.4\times 10^{-3},\quad\alpha\approx 5.2\times 10^{-3},\quad\beta\approx 0.48,\quad\eta\approx 5.4,\quad d_{0}=12.$
@@
$p^{\star}_{\mathrm{circuit}}\approx 3.1\times 10^{-3},\quad\alpha\approx 4.1\times 10^{-3},\quad\beta\approx 0.53,\quad\eta\approx 5.5,\quad d_{0}=12.$
Figure 16: Analysis of bias between $\overline{Z}$- and $\overline{X}$-failure under circuit-level noise. The left figure shows the scatter plot of the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ and the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ for all the simulation outcomes in Figs. 7 and 8. The right figure shows the scatter plot of the logical failure rate $p_{\mathrm{fail}}$ and the bias $R_{\mathrm{bias}}$ defined in Eq. (8) for all the simulation outcomes in Figs. 7 and 8. The error bars indicate the 99% confidence intervals of $R_{\mathrm{bias}}$. For both figures, the dots and error bars are colored depending on the code distance $d$ (as shown in the right color bar) and the orange dashed lines represent unbiased cases of $p_{\mathrm{fail}}^{(Z)}=p_{\mathrm{fail}}^{(X)}$. In the right figure, a distinct bias is present in a group of data points (enclosed by a dotted circle), corresponding to cases of $d=3$.
Figure 16 analyzes the bias between $\overline{Z}$- and $\overline{X}$-failure in all the circuit-level simulation outcomes of Fig. 7 and 8. It shows two scatter plots: one for the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ and the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ and the other one for the total failure rate $p_{\mathrm{fail}}$ and the bias $R_{\mathrm{bias}}$ defined in Eq. (8). We observe that the 99% confidence intervals of $R_{\mathrm{bias}}$ almost always include zero, which implies that we cannot reject the null hypothesis that there is no bias. Exceptionally, bias is not negligible when $d=3$, where $-0.31\lessapprox R_{\mathrm{bias}}\lessapprox-0.14$, meaning that $1.4\lessapprox p_{\mathrm{fail}}^{(X)}/p_{\mathrm{fail}}^{(Z)}\lessapprox 2.0$.

### v2

@@ C.2 Small-weight errors uncorrectable via the concatenated MWPM decoder
Then the subsequent MWPM on the red-only lattice produces predictions represented by black triangles in Fig. 12(b), leading to a logical error. Similarly, the green sub-decoding fails if and only if
@@
Appendix D Additional numerical analyses
@@
Figure 15 presents separate analyses for the $\overline{Z}$- and $\overline{X}$-failure of the decoder under near-threshold circuit-level noise, which are analogous to Figs. 7(b) and (c). The long-term cross thresholds $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$ are estimated as 0.460% for the $\overline{Z}$-failure and 0.452% for the $\overline{X}$-failure, which differ by about 1.8% (0.008%p). Note that $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$ for the overall failure rate is about 0.455%, which is the average of these two.
@@
$p^{\star}_{\mathrm{circuit}}\approx 3.4\times 10^{-3},\quad\alpha\approx 5.0\times 10^{-3},\quad\beta\approx 0.48,\quad\eta\approx 5.4,\quad d_{0}=12.$
@@
$p^{\star}_{\mathrm{circuit}}\approx 3.0\times 10^{-3},\quad\alpha\approx 3.4\times 10^{-3},\quad\beta\approx 0.53,\quad\eta\approx 5.5,\quad d_{0}=12.$
Figure 17: Analysis of bias between $\overline{Z}$- and $\overline{X}$-failure under circuit-level noise. The left figure shows the scatter plot of the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ and the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ for all the simulation outcomes in Figs. 7 and 8. The right figure shows the scatter plot of the logical failure rate $p_{\mathrm{fail}}$ and the bias $R_{\mathrm{bias}}$ defined in Eq. (8) for all the simulation outcomes in Figs. 7 and 8. The error bars indicate the 99% confidence intervals of $R_{\mathrm{bias}}$. For both figures, the dots and error bars are colored depending on the code distance $d$ (as shown in the right color bar) and the orange dashed lines represent unbiased cases of $p_{\mathrm{fail}}^{(Z)}=p_{\mathrm{fail}}^{(X)}$.
Figure 17 analyzes the bias between $\overline{Z}$- and $\overline{X}$-failure in all the circuit-level simulation outcomes of Fig. 7 and 8. It shows two scatter plots: one for the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ and the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ and the other one for the total failure rate $p_{\mathrm{fail}}$ and the bias $R_{\mathrm{bias}}$ defined in Eq. (8). We observe that the 99% confidence intervals of $R_{\mathrm{bias}}$ almost always include zero, which implies that we cannot reject the null hypothesis that there is no bias.